Question 4. (10 points total) (Efficiency analysis) Baryla Inc. manufactures high quality decorator lamps in a plant located in eastern Tennessee. Last year the firm had sales of $93 million and a gross profit margin of 45 percent. points) How much inventory can Baryla hold and still maintain an inventory turnover ratio of at least 6.3 times? (Round to one decimal place.) (5 points) Currently, some of Barylas inventory includes $2.3 million of outdated and damaged goods that simply remain in inventory and are not salable. What inventory ratio must the good inventory maintain in order to achieve an overall turnover ratio of at least 6.3 (including the unsalable items)? (Round to one decimal place.) b.
Answer #1

a. Sales = $93,000,000

Gross profit margin = 45%

Gross profit= 45%*93,000,000 = $41,850,000

Gross profit = sales - cost of goods sold

Cost of goods sold = Gross profit + sales = 41,850,000 + 95,000,000 = $53,150,000

Inventory turnover = cost of goods sold/inventory

Inventory = $52,250,000/6.3= $8,436,508

For inventory turnover ratio to be 6.3 times the inventory value should be $8,436,508

b. Good inventory = Total Inventory - Unsalable items = $8,436,508 - $2,300,000 = $6,136,508

The inventory turnover ratio must the good inventory maintain in order to achieve an overall turnover ratio of at least 6.3 (including the unsalable items) is

53,150,000/6,136,508 = 8.66